3y-8 y
---- = ----
12 5

Answer by checkley77(12844)

3y-8 y
---- = ----
12 5 cross multiply.
5(3y-8)=12y
15y-40=12y
15y-12y=40
3y=40
y=40/3 ans.
Proof:
(3*40/3-8)/12=(40/3)/5
(40-8)/12=40/3*1/5
32/12=40/15
32*15=12*40
480=480